OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (0, 0) and a focus at (-4, 0). PLEASE HELP! I've tried, and just can't do it..sorry.

OpenStudy (anonymous):

focus (p, 0)=(-4,0) vertex (h.k) =(0,0) x^2= (4p)*y x^2 = 4*(-4)y x^2 = -16y 1/-16x^2 = y

OpenStudy (anonymous):

wait i did it wrong

OpenStudy (anonymous):

x = -1/16y^2 

OpenStudy (anonymous):

y^2= (4p)*x y^2 = 4*(-4)x y^2 = -16x 1/-16y^2 = x

OpenStudy (anonymous):

5 bucks

OpenStudy (anonymous):

arabs lol

OpenStudy (anonymous):

yup, u?

OpenStudy (anonymous):

you know arabic?